# LancBiO: Dynamic Lanczos-aided Bilevel Optimization via Krylov Subspace

### v1

@@ I.1 General settings
We conduct experiments to empirically validate the performance of the proposed algorithms. We test on a synthetic problem, a hyper-parameters selection task, and a data hyper-cleaning task. We compare the proposed SubBiO and LancBiO with the existing algorithms in bilevel optimization: stocBiO (Ji et al., 2021), AmIGO-GD and AmIGO-CG (Arbel & Mairal, 2022), SOBA (Dagréou et al., 2022) and TTSA (Hong et al., 2023). The experiments are produced on a server that consists of two Intel® Xeon® Gold 6330 CPUs (total 2$\times$28 cores), 512GB RAM, and one NVIDIA A800 (80GB memory) GPU. The synthetic problem and the deep learning experiments are carried out on the CPUs and the GPU, respectively. For wider accessibility and application, we have made the code publicly available on https://github.com/UCAS-YanYang/LancBiO.
For the proposed LancBiO, we initiate the subspace dimension at $1$, and gradually increase it to $m=10$ for the deep learning experiments and to $m=80$ for the synthetic problem. For all the compared algorithms, we employ a grid search strategy to optimize the parameters. The optimal parameters yield the lowest loss. The experiment results are averaged over $10$ runs
@@
I.2 Synthetic problem
@@
$\begin{array}[]{cl}\min\limits_{x\in\mathbb{R}^{d}}&f(x,y):=c_{1}\cos\left({x^{\top}D_{1}y}\right)+\frac{1}{2}\left\|{D_{2}x-y}\right\|^{2},\\[14.22636pt] \mathrm{s.\,t.}&\begin{aligned} y^{*}&=\operatorname*{arg\,min}\limits_{y\in\mathbb{R}^{d}}g(x,y)\\ :&=c_{2}\sum_{i=1}^{d}{\sin(x_{i}+y_{i})}+\log\left({\sum_{i=1}^{d}{e^{x_{i}y_{i}}}}\right)+\frac{1}{2}y^{\top}\left({D_{3}+G}\right)y,\end{aligned}\end{array}$
@@
When $m=50$, the estimation of $v^{*}$ is sufficiently accurate to facilitate effective hyper-gradient convergence, which is demonstrated in Figure 5 that for $m\geq 50$, further increases in $m$ merely enhance the convergence of the residual norm. Under the same outer iterations, to attain a comparable convergence property, $I$ for AmIGO-CG should be set to $10$. Furthermore, given that the number of Hessian-vector products averages at $(1+\frac{1}{m})$ per outer iteration for LancBiO, whereas AmIGO requires $I\geq 2$, it follows that LancBiO is more efficient.
I.3 Data hyper-cleaning on MNIST
The data hyper-cleaning task (Shaban et al., 2019), conducted on the MNIST dataset (LeCun et al., 1998), aims to train a classifier in a corruption scenario, where the labels of the training data are randomly altered to incorrect classification numbers at a certain probability $p$, referred to as the corruption rate. The task is formulated as follows,
In the deterministic setting, where we implement all compared methods with full-batch, the training set, the validation set and the test set contain $5000$, $5000$ and $10000$ samples, respectively. For algorithms that incorporate inner iterations to approximate $y^{*}$ or $v^{*}$, we select the inner iteration number from the set $\left\{5i\mid i=1,2,3,4\right\}$. The step size of inner iteration is selected from the set $\{0.01,0.1,1,10\}$ and the step size of outer iteration is chosen from $\left\{5\times 10^{i}\mid i=-3,-2,-1,0,1,2,3\right\}$. The results are presented in Figure 6. Note that LancBiO is crafted for approximating the Hessian inverse vector product $v^{*}$, while the solid methods stocBiO and TTSA are not. Consequently, with respect to the residual norm of the linear system, i.e., $\left\|{A_{k}v_{k}-b_{k}}\right\|$, we only compare the results with AmIGO-GD, AmIGO-CG and SOBA. Observe that the proposed subspace-based LancBiO achieves the lowest residual norm and the best test accuracy, and subBiO is comparable to the other algorithms. Specifically, in Figure 6, the efficiency of LancBiO stems from its accurate approximation of the linear system. Furthermore, we implement the solvers designed for the stochastic setting using mini-batch to enable a broader comparison in Figure 9. It is shown that the stochastic algorithm SOBA tends to converge faster initially, but algorithms employing a full-batch approach achieve higher accuracy.
Generally, for standard linear systems, the traditional Lanczos process is recognized for its efficiency and versatility over gradient descent methods. LancBiO, in a sense, reflects this principle within the context of bilevel optimization, underscoring the effectiveness of the dynamic Lanczos-aided approach.
@@ I.4 Logistic regression on $20$Newsgroup
Consider the hyper-parameters selection task on the $20$Newsgroups dataset (Grazzi et al., 2020), which contains $c=20$ topics with around $18000$ newsgroups posts represented in a feature space of dimension $l=130107$. The goal is to simultaneously train a linear classifier $w$ and determine the optimal regularization parameter $\zeta$. The task is formulated as follows,
@@
Figure 11: Comparison of the bilevel algorithms hyper-parameters selection task. Left: validation accuracy; Center: validation loss; Right: residual norm of the linear system, $\left\|{A_{k}v_{k}-b_{k}}\right\|$.

### v2

@@ I.1 General settings
We conduct experiments to empirically validate the performance of the proposed algorithms. We test on a synthetic problem, a hyper-parameters selection task, and a data hyper-cleaning task. We compare the proposed SubBiO and LancBiO with the existing algorithms in bilevel optimization: stocBiO [Ji et al., 2021], AmIGO-GD and AmIGO-CG [Arbel and Mairal, 2022], SOBA [Dagréou et al., 2022] and TTSA [Hong et al., 2023], F2SA [Kwon et al., 2023] and HJFBiO [Huang, 2024]. The experiments are produced on a workstation that consists of two Intel® Xeon® Gold 6330 CPUs (total 2$\times$28 cores), 512GB RAM, and one NVIDIA A800 (80GB memory) GPU. The synthetic problem and the deep learning experiments are carried out on the CPUs and the GPU, respectively. For wider accessibility and application, we have made the code available on https://github.com/UCAS-YanYang/LancBiO.
For the proposed LancBiO, we initiate the subspace dimension at $1$, and gradually increase it to $m=10$ for the deep learning experiments and to $m=80$ for the synthetic problem. For all the compared algorithms, we employ a grid search strategy to optimize the parameters. The optimal parameters yield the lowest loss. The experiment results are averaged over $10$ runs. Note that 3.9 for initialization is not used in practice, for which we treat it as a theoretical assumption rather than incorporating it into Algorithm 2 in this paper.
@@
I.2 Data hyper-cleaning
The data hyper-cleaning task [Shaban et al., 2019], conducted on the MNIST dataset [LeCun et al., 1998], aims to train a classifier in a corruption scenario, where the labels of the training data are randomly altered to incorrect classification numbers at a certain probability $p$, referred to as the corruption rate. The task is formulated as follows,
In the deterministic setting, where we implement all compared methods with full-batch, the training set, the validation set and the test set contain $5000$, $5000$ and $10000$ samples, respectively. For algorithms that incorporate inner iterations to approximate $y^{*}$ or $v^{*}$, we select the inner iteration number from the set $\left\{5i\mid i=1,2,3,4\right\}$. The step size of inner iteration is selected from the set $\{0.01,0.1,1,10\}$ and the step size of outer iteration is chosen from $\left\{5\times 10^{i}\mid i=-3,-2,-1,0,1,2,3\right\}$. Regarding the Hessian/Jacobian-free algorithm HJFBiO, we set the step size $\delta=1\times 10^{-5}$ to implement finite difference methods. The results are presented in Figure 4. Note that LancBiO is crafted for approximating the Hessian inverse vector product $v^{*}$, while the solid methods stocBiO, TTSA, F2SA, and HJFBiO are not. Consequently, concerning the residual norm of the linear system, i.e., $\left\|{A_{k}v_{k}-b_{k}}\right\|$, we only compare the results with AmIGO-GD, AmIGO-CG and SOBA. Observe that the proposed subspace-based LancBiO achieves the lowest residual norm and the best test accuracy, and SubBiO is comparable to the other algorithms. Specifically, in Figure 4, the efficiency of LancBiO stems from its accurate approximation of the linear system. Furthermore, we implement the solvers designed for the stochastic setting using mini-batch to enable a broader comparison in Figure 8. It is shown that the stochastic algorithm SOBA tends to converge faster initially, but algorithms employing a full-batch approach achieve higher accuracy.
Additionally, we also evaluate the performance of bilevel algorithms on Fashion-MNIST [Xiao et al., 2017] and Kuzushiji-MNIST [Clanuwat et al., 2018] datasets, both of which present more complexity compared to MNIST. Specifically, Fashion-MNIST serves as a modern replacement for MNIST, featuring grayscale images of clothing items across $10$ categories, and Kuzushiji-MNIST is a culturally rich dataset of handwritten Japanese characters. The results, reported in Figures 10 and 11, reveal that LancBiO performs better than other algorithms and showcases robustness across various datasets, and SubBiO delivers a comparable convergence property.
Figure 10: Data hyper-cleaning task tested on the Fashion-MNIST dataset when $p=0.5$. Left: test accuracy; Center: test loss; Right: residual norm of the linear system, $\left\|{A_{k}v_{k}-b_{k}}\right\|$.
Figure 11: Data hyper-cleaning task tested on the Kuzushiji-MNIST dataset when $p=0.6$. Left: test accuracy; Center: test loss; Right: residual norm of the linear system, $\left\|{A_{k}v_{k}-b_{k}}\right\|$.
I.3 Synthetic problems
@@
$\begin{array}[]{cl}\min\limits_{x\in\mathbb{R}^{d}}&f(x,y^{*}):=c_{1}\cos\left({x^{\top}D_{1}y^{*}}\right)+\frac{1}{2}\left\|{D_{2}x-y^{*}}\right\|^{2},\\[5.69054pt] \mathrm{s.\,t.}&\begin{aligned} y^{*}&=\operatorname*{arg\,min}\limits_{y\in\mathbb{R}^{d}}g(x,y)\\ :&=c_{2}\sum_{i=1}^{d}{\sin(x_{i}+y_{i})}+\log\left({\sum_{i=1}^{d}{e^{x_{i}y_{i}}}}\right)+\frac{1}{2}y^{\top}\left({D_{3}+G}\right)y,\end{aligned}\end{array}$ (72)
@@
When $m=50$, the estimation of $v^{*}$ is sufficiently accurate to facilitate effective hyper-gradient convergence, which is demonstrated in Figure 5 that for $m\geq 50$, further increases in $m$ merely enhance the convergence of the residual norm. Under the same outer iterations, to attain a comparable convergence property, $I$ for AmIGO-CG should be set to $10$. Furthermore, given that the number of Hessian-vector products averages at $(1+{1}/{m})$ per outer iteration for LancBiO, whereas AmIGO requires $I\geq 2$, it follows that LancBiO is more efficient.
Table 2: Comparison on the synthetic problem (72). The dimension of problems is denoted by $d$. Results are averaged over 10 runs.
Algorithm $d=10$ $d=100$ $d=1000$ $d=10000$
UL Val. Time (S) UL Val. Time (S) UL Val. Time (S) UL Val. Time (S)
LancBiO ${4.52}\mathrm{e}{-2}$ 0.320.32{0.32}0.32 ${6.37}\mathrm{e}{-2}$ $0.53$ ${5.29}\mathrm{e}{-2}$ $1.30$ ${-1.21}\mathrm{e}{-2}$ 16.3616.36\ \ \ 16.3616.36
SubBiO ${3.73}\mathrm{e}{-2}$ 1.001.00{1.00}1.00 ${7.19}\mathrm{e}{-2}$ $1.17$ ${4.63}\mathrm{e}{-2}$ $1.72$ ${-2.91}\mathrm{e}{-2}$ 21.2621.26\ \ \ 21.2621.26
AmIGO-GD ${1.67}\mathrm{e}{-1}$ 2.442.44{2.44}2.44 ${1.05}\mathrm{e}{-1}$ $3.48$ ${1.05}\mathrm{e}{-1}$ $1.46$ $\ \ \ {4.96}\mathrm{e}{-2}$ 46.5346.53\ \ \ 46.5346.53
AmIGO-CG ${5.56}\mathrm{e}{-2}$ 0.400.40{0.40}0.40 ${7.65}\mathrm{e}{-2}$ $1.90$ ${5.73}\mathrm{e}{-2}$ $2.44$ $\ \ \ {3.68}\mathrm{e}{-2}$ 25.0025.00\ \ \ 25.0025.00
SOBA ${1.70}\mathrm{e}{-1}$ $0.60$ ${1.28}\mathrm{e}{-1}$ $1.64$ ${1.03}\mathrm{e}{-1}$ $2.52$ ${\ \ \ 3.49}\mathrm{e}{-2}$ 33.8933.89{\ \ \ 33.89}33.89
TTSA ${5.66}\mathrm{e}{-2}$ 0.470.47{0.47}0.47 ${5.52}\mathrm{e}{-2}$ $0.89$ ${6.52}\mathrm{e}{-2}$ $2.81$ ${\ \ \ 1.87}\mathrm{e}{-1}$ $121.07$
stocBiO ${6.24}\mathrm{e}{-2}$ 0.290.29{0.29}0.29 ${6.02}\mathrm{e}{-2}$ $0.45$ ${5.21}\mathrm{e}{-2}$ $1.34$ ${-1.30}\mathrm{e}{-2}$ 20.6620.66\ \ \ 20.6620.66
Moreover, to illustrate how the proposed methods scale with increasing dimensions, we present the convergence time and the final upper-level (UL) value under different problem dimensions $d=10^{i},i=1,2,3,4$ in Table 2. The results show the robustness of the proposed methods across varying problem dimensions.
In addition, as discussed in Appendix D, to address the bilevel problem where the lower-level problem exhibits an indefinite Hessian, the framework LancBiO (Algorithm 2) requires a minor modification. Specifically, line 13 in Algorithm 2, which solves a small-size tridiagonal linear system, will be replaced by solving a low-dimensional least squares problem. We test the modified method LancBiO-MINRES on the following bilevel example borrowed from Liu et al. [2023a] with a non-convex lower-level problem,
$\begin{array}[]{cl}\min\limits_{x\in\mathbb{R}}&f(x,y^{*}):=\left\|{x-a}\right\|^{2}+\left\|{y^{*}-a-c}\right\|^{2}\\[5.69054pt] \mathrm{s.\,t.}&y_{i}^{*}\in\operatorname*{arg\,min}\limits_{y_{i}\in\mathbb{R}}\,\sin(x+y_{i}-c_{i}),\,\text{for}\,i=1,2,\ldots,d,\end{array}$ (73)
where the subscript $i$ denotes the $i$-th component of a vector, while $a\in\mathbb{R}$ and $c\in\mathbb{R}^{d}$ are parameters. The results, reported in Figure 13, imply the potential of extending our work to non-convex scenarios.
Figure 13: Test LancBiO-MINRES on the synthetic problem (73) with $d=100$. The metric follows from the necessary conditions developed for bilevel problems without lower-level convexity in Theorem 1 of Xiao et al. [2023].
@@ I.4 Logistic regression on $20$Newsgroup
Consider the hyper-parameters selection task on the $20$Newsgroups dataset [Grazzi et al., 2020], which contains $c=20$ topics with around $18000$ newsgroups posts represented in a feature space of dimension $l=130107$. The goal is to simultaneously train a linear classifier $w$ and determine the optimal regularization parameter $\zeta$. The task is formulated as follows,